OpenStudy (anonymous):

You have 80 yards of fencing to enclose a rectangular region.find the dimensions of the rectangle that maximize the enclose area.what is the maximize area? I NEED ALL THE STEPS!

OpenStudy (raden):

actually, if u want finish it soon u can suppose the rectangle has to be a square so, what's the length of side a square when its perimeter = 80, obviously, 20 right ? so, the area = side * side = 20 * 20 = ....

OpenStudy (anonymous):

P=80=2x + 2y your area formula is A=xy Since you have your P equation... you can get your x in terms of y by \[\frac{ 2x+2y=80 }{ 2 }\] which thne you get\[x+y=40\] so then you get \[x=40-y\] plug in x for your equation in your area so you get... \[A = (40-y)y\] Distribute... so you get \[A = 40y - y^{2}\] take the derivative of A so that you get.. \[A' = 40 - 2y\] set A' = 0 and solve... \[0 = 40-2y\] \[-40 = -2y\] \[20 = y\] plug back in y for your x equation \[x = 40-y\] \[x = 40-20\] \[x = 20\]

OpenStudy (raden):

the finally, the pic be a square :p

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

yeah it is...

OpenStudy (anonymous):

thanks^-^

OpenStudy (anonymous):

wait so the answer is 400 then right?

OpenStudy (anonymous):

yupp